OpenStudy (anonymous):

Having a little trouble with it: Is \(f(x) = \dfrac{2x}{1+x^2}\) an injective? Why or why not? If I graph this function, we can see that it is not injective, but how to show that algebraically?

OpenStudy (anonymous):

@ganeshie8 @iGreen @AriPotta

OpenStudy (anonymous):

OK, we start with \(f(a) = f(b)\) I am stuck at \(b+a^2b = a+ab^2\)... I honestly don't know how to show that it is not injective without depending on graph.

OpenStudy (zarkon):

what is \[f(1/x)\]

OpenStudy (uri):

We can go from there: